### v2

# Compatibility of Kazhdan and Brauer Homomorphism

Abstract.
Let $G$ be a split connected reductive group, defined over $\mathbb{Z}$. Let $F$ and $F^{\prime}$ be the non-archimedean $m$-close local fields, where $m$ is a positive integer. D.Kazhdan gave an isomorphism between the Hecke algebras ${\rm Kaz}_{m}^{F}:\mathcal{H}\big{(}G(F),K_{F}\big{)}\rightarrow\mathcal{H}\big{(}G(F^{\prime}),K_{F^{\prime}}\big{)}$, where $K_{F}$ and $K_{F^{\prime}}$ are the $m$-th usual congruence subgroups of $G(F)$ and $G(F^{\prime})$ respectively. On the other hand, if $\sigma$ is an automorphism of $G$ of prime order $l$, then we have Brauer homomorphism ${\rm Br}:\mathcal{H}(G(F),U(F))\rightarrow\mathcal{H}(G^{\sigma}(F),U^{\sigma}(F))$, where $U(F)$ and $U^{\sigma}(F)$ are compact open subgroups of $G(F)$ and $G^{\sigma}(F)$ respectively. This article aims to study the compatibility between these two maps in certain sense. Further, an application of this compatibility is given in the context of linkage–which is the representation theoretic version of Brauer homomorphism.
1. Introduction
The local Langlands correspondence relates the set of irreducible smooth representations of the group of rational points of a reductive group over a local field–with the representation theory of its Weil–Deligne group. For a split reductive group G defined over $\mathbb{Z}$, Kazhdan conjectured a link between the local Langlands correspondences for $\operatorname{{\bf G}}(F)$ and $\operatorname{{\bf G}}(F^{\prime})$, where $F$ is a $p$-adic field and $F^{\prime}$ is a non-Archimedean equal characteristic local field which are sufficiently close. Kazhdan’s approach is via isomorphisms between the respective Hecke algebras, with integral coefficients, $\mathcal{H}(\operatorname{{\bf G}}(F),K)$ and $\mathcal{H}(\operatorname{{\bf G}}(F^{\prime}),K^{\prime})$, for some specific choice of open-compact subgroups $K$ and $K^{\prime}$ of $\operatorname{{\bf G}}(F)$ and $\operatorname{{\bf G}}(F^{\prime})$ respectively (see [Kaz86, Theorem A]). Say $L$ is a non-Archimedean local field with residue characteristic $p$ and assume that $K$ is a nice compact open subgroup of $\operatorname{{\bf G}}(L)$–for instance a congruence subgroup of positive level. When $\operatorname{{\bf G}}(L)$ has an automorphism, denoted by $\sigma$, of prime order $l\neq p$ such that $\sigma(K)=K$, Treumann–Venkatesh ([TV16, Section 4]) define an $\overline{\mathbb{F}_{l}}$-algebra homomorphism between $\mathcal{H}(\operatorname{{\bf G}}(L),K)\otimes\overline{\mathbb{F}_{l}}$ and $\mathcal{H}(\operatorname{{\bf G}}^{\sigma}(L),K^{\sigma})\otimes\overline{\mathbb{F}_{l}}$, which is called the Brauer homomorphism. The Brauer homomorphism is conjectured to be compatible with Langlands functoriality principle. In this article, we study the compatibility of Brauer homomorhism and Kazhdan homomorphism in the Base change setting. Let $\mathfrak{o}_{L}$ be the ring of integers of $L$ and let $\mathfrak{p}_{L}$ be the maximal ideal of $\mathfrak{o}_{L}$. Let $E$ be a finite extension of $L$ of prime degree $l$, and let $\sigma$ be the generator of ${\rm Gal}(E/L)$. Let $K_{E}$ be the $m$-th congruence subgroup of $\operatorname{{\bf G}}(\mathfrak{o}_{E})$. Note that $K_{E}$ is stable under the action of $\sigma$ and let $K_{L}=K_{E}\cap\operatorname{{\bf G}}(L)$. Truemann and Venkatesh define Brauer homomorphism ${\rm Br}$, which is the algebra homomorphism
${\rm Br}:\mathcal{H}(\operatorname{{\bf G}}(E),K_{E})^{\sigma}\rightarrow\mathcal{H}(\operatorname{{\bf G}}(L),K_{L})$
obtained by restriction of functions on $\operatorname{{\bf G}}(E)$ to $\operatorname{{\bf G}}(L)$(see [TV16, Section 4]). The homomorphism ${\rm Br}$ is conjectured to be compatible with Langlands functoriality principle. Kazhdan ([Kaz86, Theorem A]) showed that there exists an integer $r\geq m$ such that if $F$ and $F^{\prime}$ are $r$-close (see Section (4)), there is a Hecke algebra isomorphism
${\rm Kaz}_{m}^{F}:\mathcal{H}(\operatorname{{\bf G}}(F),K_{F})\rightarrow\mathcal{H}(\operatorname{{\bf G}}(F^{\prime}),K_{F^{\prime}}).$
Kazhdan conjectured that it is sufficient to take $r=m$, and this conjecture is proved by Ganapathy in [Gan15]. It is shown in [Gan15] that Kazhdan isomorphism is compatible with the various properties in representation theory, and it also establish a bridge between local Langlands correspondence(LLC) over the field characteristic zero and that of positive characteristic.
We now set up some notations to formulate the main result of this article. Let $E$ be a finite Galois extension of a non-archimedean local field $F$ of prime degree $l$. Let $F^{\prime}$ be another non-Archimedean local field which is $m$-close to $F$. Throughout this article, we assume that the prime number $l$ is different from $p$, the residue characteristic of both $F$ and $F^{\prime}$. Lemma 4.2 in Section 4 says that we have a finite Galois extension $E^{\prime}$ of $F^{\prime}$ of degree $l$ such that the fields $E$ and $E^{\prime}$ are $em$-close, where $e=1$ if $E$ is unramified and $e=l$ if $E$ is totally ramified. Let $\sigma$ (resp. $\sigma^{\prime}$) be a generator of the cyclic group ${\rm Gal}(E/F)$ (resp. ${\rm Gal}(E^{\prime}/F^{\prime})$). This leads to the following diagram:
[…]
In this article, we address the natural question about the commutativity of the above diagram. Let us state it as a theorem.
Theorem 1.1.
Let $F$ and $F^{\prime}$ be the non-archimedean $m$-close local fields. Let $E$ be the finite Galois extensions of $F$ of prime degree $l$, and let $E^{\prime}$ be the Galois extension of $F^{\prime}$, as indicated above. Then for any split connected reductive group G defined over $\mathbb{Z}$, we have
${\rm Br}^{\prime}\circ{\rm Kaz}_{em}^{E}={\rm Kaz}_{m}^{F}\circ{\rm Br}.$
In the context of representation theory, there is a notion of linkage (see subsection 6.2, for definition), defined by Treumann-Venkatesh–which is the representation theoretic analouge of Brauer homomorphism in the sense that they are compatible ([TV16, Section 6]). In this article, we give a precise formulation (subsection 6.3) of the compatibility of Brauer homomorphism with linkage. Then, we use this formulation and Theorem 1.1 to establish the compatibility of linkage with Kazhdan isomorphism (Theorem 6.3).
Now we briefly explain the contents of this article. In Section $2$, we review the Hecke algebra structure associated to any locally profinite group. In Section $3$, we formulate the Brauer homomorphism in our setting. The Kazdhan isomorphism and its compatibility with Galois actions will be discussed in Section $4$. In Section $5$, we prove Theorem 1.1. In Section $6$, we recall the notion of linkage and formulate some finiteness results on Tate cohomology group. Following this, we discuss the compatibility of linkage with Kazhdan isomorphism.
2. Hecke Algebra
In this section, we review the Hecke algebra structure on any locally profinite group. We refer to [TV16, Subsection 2.10] for more insight.
2.1.
Let $k$ be an algebraically closed field of positive characteristic $l$. Let $G$ be a locally profinite (i.e., locally compact and totally disconnected) group. Let $K$ be a compact open subgroup of $G$. We denote by $G/K$ the set of all distinct left cosets of $K$ in $G$, and it is a discrete set with a left action of $G$. For any $g\in G$, the left coset $gK$ is sometimes denoted by the symbol $[g]$. There is a natural action of $G\times G$, via left translation, on the space of all $k$-valued functions on the discrete set $G/K\times G/K$. We denote by $\mathcal{F}(G//K)$ the space of all such $k$-valued functions on $G/K\times G/K$ that are invariant under the action of the diagonal subgroup $\Delta G=\{(g,g):g\in G\}$, and whose support is a union of finitely many $G$-orbits. There is a multiplicative structure on the space $\mathcal{F}(G//K)$, given by
$(f_{1}*f_{2})([x],[z])=\sum_{y\in G/K}f_{1}([x],[y])f_{2}([y],[z]),$
for all $[x],[z]\in G/K$ and $f_{1},f_{2}\in\mathcal{F}(G//K)$. The space $\mathcal{F}(G//K)$ with the above multiplicative structure is a $k$-algebra, and it is called the Hecke algebra of $G$ associated with $K$. We now give an equivalent description of the Hecke algebra, which is relevant for our context.
2.2.
Let us consider the equivalence relation $\sim$ on the set $G/K\times G/K$, where
$([x],[y])\sim([x^{\prime}],[y^{\prime}])$
if and only if there exists $g\in G$ such that $[gx]=[gx^{\prime}]$ and $[gy]=[gy^{\prime}]$. Then for any $[g],[h]\in G/K$, the map $([g],[h])\mapsto Kg^{-1}hK$ induces a bijection from the set of equivalence classes $(G/K\times G/K)/\sim$ onto the double coset space $K\setminus G/K$.
Let $\mathcal{H}(G,K)$ be the space of all $k$-valued functions which are bi-$K$-invariant i.e., $f(k_{1}gk_{2})=f(g),$ for all $k_{1},k_{2}\in K$, $g\in G$ and $f\in\mathcal{H}(G,K)$. Then the preceding bijection gives the following map
$\mathcal{F}(G//K)\longrightarrow\mathcal{H}(G,K)$
$\phi\longmapsto\widetilde{\phi},$ (2.1)
where the function $\widetilde{\phi}:G\rightarrow k$ is defined as $\widetilde{\phi}(g):=\phi([1],[g])$, for all $g\in G$. The map (2.1) is clearly a bijection, and this induces a $k$-algebra structure on the space $\mathcal{H}(G,K)$. In the literature, many authors refer to the $k$-algebra $\mathcal{H}(G,K)$ as the Hecke algebra of $G$ associated with $K$.
3. Brauer Homomorphism
3.1.
Let $G$ be a locally profinite group, and let $\sigma$ be an automorphism of $G$ of order $l$. Let $K$ be a compact open subgroup of $G$ with $\sigma(K)=K$. Let $G^{\sigma}$ (resp. $K^{\sigma}$) be the the fixed subgroup of $G$ (resp. $K$) under the action of $\sigma$. Moreover, the compact open subgroup $K$ is called $\sigma$-plain if
  1. (1)
    The inclusion
    $G^{\sigma}/K^{\sigma}\hookrightarrow(G/K)^{\sigma}$
    $gK^{\sigma}\mapsto gK$
    is a bijection.
  2. (2)
    There exists a subgroup $K^{\prime}$ of $K$ of finite index such that $K^{\prime}$ is an inverse limit of finite subgroups each of which is coprime to $l$.
3.2.
For any $\sigma$-plain compact open subgroup $K$ of $G$, the discrete set $G/K\times G/K$ is equipped with an action of the group $\langle\sigma\rangle$, defined as
$\sigma.([g],[h]):=\big{(}[\sigma(g)],[\sigma(h)]\big{)},$
for all $[g],[h]\in G/K$. This action further factorizes through the quotient space $G/K\times G/K/\sim$, and it is given by
$\sigma.\overline{([g],[h])}=\overline{([\sigma(g)],[\sigma(h)])},$
where $\overline{([g],[h])}$ denotes the equivalence class of the element $([g],[h])$. The space $\mathcal{F}(G//K)$ carries a natural action of $\langle\sigma\rangle$, defined by
$(\sigma.\phi)([x],[y])=\phi\big{(}[\sigma^{-1}(x)],[\sigma^{-1}(y)]\big{)},$
for all $x,y\in G$ and $\phi\in\mathcal{F}(G//K)$. It is shown in [TV16, Section 4] that the Brauer homomorphism ${\rm Br}$ is a $k$-algebra homomorphism from $\mathcal{F}(G//K)^{\sigma}$ to $\mathcal{F}(G^{\sigma}//K^{\sigma})$, and it is the restriction map
${\rm Br}(\phi):={\rm Res}_{(G^{\sigma}/K^{\sigma}\times G^{\sigma}/K^{\sigma})}(\phi)$ (3.1)
for all $\phi\in\mathcal{F}(G//K)^{\sigma}$.
3.3.
Now, we formulate the Brauer homomorphism (3.1) for the $k$-algebra $\mathcal{H}(G,K)$, defined as in the preceding section. First, let us consider the action of the group $\langle\sigma\rangle$ on the double coset space $K\setminus G/K$ :
$\sigma.(KgK):=K\sigma(g)K,$
for all $g\in G$. Then we have a natural induced action of $\langle\sigma\rangle$ on the space $\mathcal{H}(G,K)$, defined as
$(\sigma.f)(g):=f(\sigma^{-1}(g)),$
for all $g\in G$. Note that the bijection (2.1) preserves the action of the group $\langle\sigma\rangle$. Then we have commutative diagram
${\mathcal{F}(G//K)^{\sigma}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\rm Br}}$${\mathcal{H}(G,K)^{\sigma}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\rm\overline{Br}}}$${\mathcal{F}(G^{\sigma}//K^{\sigma})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{H}(G^{\sigma},K^{\sigma})}$
where the horizontal arrows in the diagram are the bijections induced by (2.1). We observe that the $k$-algebra map ${\rm\overline{Br}}:\mathcal{H}(G,K)^{\sigma}\rightarrow\mathcal{H}(G^{\sigma},K^{\sigma})$, induced by the above diagram, is the restriction map to subgroup $G^{\sigma}$ i.e.,
${\rm\overline{Br}}(f):={\rm Res}_{G^{\sigma}}(f),$
for all $f\in\mathcal{H}(G,K)^{\sigma}$
4. Kazhdan isomorphism
In this section, we review Kazhdan isomorphism and set up some initial results. Let $F$ and $F^{\prime}$ be the non-Archimedean local fields, both of residue characteristic $p$, with ring of integers $\mathfrak{o}_{F}$ and $\mathfrak{o}_{F^{\prime}}$ and its maximal ideals $\mathfrak{p}_{F}$ and $\mathfrak{p}_{F^{\prime}}$ respectively. The fields $F$ and $F^{\prime}$ are called $m$-close if there is a ring isomorphism $\Lambda:\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}\rightarrow\mathfrak{o}_{F^{\prime}}/\mathfrak{p}_{F^{\prime}}^{m}$. We choose uniformizer $\varpi$ (resp. $\varpi^{\prime}$) of the field $F$ (resp. $F^{\prime}$) such that the class of $\varpi$ corresponds to the class of $\varpi^{\prime}$ via the map $\Lambda$.
4.1.
Let G be a split connected reductive group defined over $\mathbb{Z}$. Fix a Borel subgroup B of G, defined over $\mathbb{Z}$. Let $\textbf{B}=\textbf{T}\textbf{U}$ be a Levi decomposition over $\mathbb{Z}$ with maximal torus T and unipotent radical U. We denote by $X^{*}(\textbf{T})$ and $X_{*}(\textbf{T})$ the character lattice and the cocharacter lattice of T respectively. Then there is a natural $\mathbb{Z}$-bilinear pairing
$X^{*}(\textbf{T})\times X_{*}(\textbf{T})\longrightarrow\mathbb{Z}$
$(\alpha,\lambda)\longmapsto\langle\alpha,\lambda\rangle,$
where $(\lambda\circ\alpha)(x)=x^{\langle\alpha,\lambda\rangle}$, for all $x\in\mathbb{G}_{m}$. Let $\Phi\subseteq X^{*}(\textbf{T})$ be the set of roots of T in G, and let $\Phi^{+}$ be the set of positive roots of T in B. Let $G_{F}=\textbf{G}(F)$. Consider the subgroup
$K_{F}={\rm Ker}\big{(}\textbf{G}(\mathfrak{o}_{F})\longrightarrow\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})\big{)}.$
Then $K_{F}$ is a compact open subgroup of $G_{F}$. Fix a Haar measure $\mu_{F}$ on $G_{F}$ such that $\mu_{F}(K_{F})=1$. For $g\in G_{F}$, let $t_{g}=1_{K_{F}gK_{F}}$, the characteristic function on the double coset $K_{F}gK_{F}$. Then the Hecke algebra $\mathcal{H}(G_{F},K_{F})$ is generated as a $k$-vector space by $\{t_{g}:g\in G_{F}\}$. We denote by $\mathcal{H}_{\mathbb{Z}}(G_{F},K_{F})$ the space generated by $\{t_{g}:g\in G_{F}\}$ over $\mathbb{Z}$. Set
$X_{*}(\textbf{T})^{-}=\{\lambda\in X_{*}(\textbf{T}):\langle\alpha,\lambda\rangle\leq 0,\alpha\in\Phi^{+}\}.$
We have the Cartan decomposition
$G_{F}=\coprod_{\mu\in X_{*}(\textbf{T})^{-}}\textbf{G}(\mathfrak{o}_{F})\varpi_{\mu}\textbf{G}(\mathfrak{o}_{F}),$ (4.1)
where $\varpi_{\mu}=\mu(\varpi)$. For each $\mu\in X_{*}(\textbf{T})^{-}$, the double coset $G_{F}(\mu):=\textbf{G}(\mathfrak{o}_{F})\varpi_{\mu}\textbf{G}(\mathfrak{o}_{F})$ is a homogeneous space under the action of the group $\textbf{G}(\mathfrak{o}_{F})\times\textbf{G}(\mathfrak{o}_{F})$, defined as
$(x,y).g:=xgy^{-1},$
for all $x,y\in\textbf{G}(\mathfrak{o}_{F})$ and $g\in G_{F}(\mu)$. Then the finite group $H_{m}:=\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})\times\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})$ acts transitively on the set $\{K_{F}xK_{F}:x\in G_{F}(\mu)\}$. Let $\Gamma_{\mu}(F)$ be the stabilizer of $K_{F}\varpi_{\mu}K_{F}$ in $H_{m}$. Let $T_{\mu}(F)\subseteq\textbf{G}(\mathfrak{o}_{F})\times\textbf{G}(\mathfrak{o}_{F})$ be the set of representatives of the coset space $H_{m}/\Gamma_{\mu}(F)$ under the composition
$\textbf{G}(\mathfrak{o}_{F})\times\textbf{G}(\mathfrak{o}_{F})\xrightarrow{{\rm mod}-\mathfrak{p}_{F}^{m}}H_{m}\longrightarrow H_{m}/\Gamma_{\mu}(F)$
Then (4.1) can be re-written as
$G_{F}=\coprod_{\mu\in X_{*}(\textbf{T})^{-}}\coprod_{(b_{i},b_{j})\in T_{\mu}(F)}K_{F}b_{i}\varpi_{\mu}b_{j}^{-1}K_{F}.$ (4.2)
For each $\mu\in X_{*}(\textbf{T})^{-}$, the map $\Lambda$ induces a bijection $T_{\mu}(F)\rightarrow T_{\mu}(F^{\prime})$. Under this bijection, there is an isomorphism ([Kaz86, Section 1]) of $\mathbb{Z}$-modules:
$\mathcal{H}_{\mathbb{Z}}(G_{F},K_{F})\longrightarrow\mathcal{H}_{\mathbb{Z}}(G_{F^{\prime}},K_{F^{\prime}})$
$t_{a_{i}\varpi_{\lambda}a_{j}^{-1}}\longmapsto t_{a_{i}^{\prime}\varpi_{\lambda}^{\prime}a_{j}^{\prime-1}},$
and this induces a vector space isomorphism
${\rm Kaz}_{m}^{F}:\mathcal{H}(G_{F},K_{F})\longrightarrow\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}}),$
where $\lambda\in X_{*}(\textbf{T})^{-}$, $(a_{i},a_{j})\in T_{\lambda}(F)$ and $(a_{i}^{\prime},a_{j}^{\prime})$ is the corresponding element in $T_{\lambda}(F^{\prime})$. For $F$ and $F^{\prime}$, sufficiently close fields, the following result is due to Kazhdan ([Kaz86, Theorem A]).
Proposition 4.1.
Given $m\geq 1$. There exists a positive integer $r\geq m$ such that if $F$ and $F^{\prime}$ are $r$-close, the map ${\rm Kaz}_{m}^{F}$ defined above is an algebra isomorphism.
Kazhdan conjectured that one could take $r=m$ in Proposition 4.1, as later proved by R.Ganapathy in her paper [Gan15].
4.2.
Let $F$ and $F^{\prime}$ be the non-Archimedean $m$-close local fields with residue chracteristic $p$. In [Del84], Deligne proved that there is an isomorphism
${\rm Gal}(\overline{F}/F)/\mathcal{I}_{F}^{m}\xrightarrow{{\rm Del}_{m}}{\rm Gal}(\overline{F^{\prime}}/F^{\prime})/\mathcal{I}_{F^{\prime}}^{m},$
where $\overline{F}$ (resp. $\overline{F^{\prime}}$) is the seperable closure of $F$ (resp. $F^{\prime}$), and $\mathcal{I}_{F}^{m}$ (resp. $\mathcal{I}_{F^{\prime}}^{m}$) is the $m$-th higher ramification subgroup of the inertia group $\mathcal{I}_{F}$ (resp. $\mathcal{I}_{F^{\prime}}$). Let $E$ be a finite Galois extension of $F$ of prime degree $l$ with $l\neq p$. Following the description of the map ${\rm Del}_{m}$, as mentioned in [Gan15, Section 2.1], we have a finite Galois extension $E^{\prime}$ of $F^{\prime}$ of same degree $l$. Moreover, the extension $E^{\prime}/F^{\prime}$ can be constructed so that $E$ and $E^{\prime}$ are likewise sufficiently close, as ensured by the next lemma.
Lemma 4.2.
Let $F$ and $F^{\prime}$ be the $m$-close non-Archimedean local fields, and let $E$ be the finite Galois extension of $F$ of prime degree $l$. Then there exists a finite Galois extension $E^{\prime}$ of $F^{\prime}$ such that $E$ and $E^{\prime}$ are $em$-close, where $e\in\{1,l\}$.
Proof.
The proof is divided into two parts. In the first part, we consider the case when $E$ is unramified over $F$. The second part deals with totally ramified case. Note that the extension $E/F$ is tamely ramified as $l\neq p$. We denote by $A_{m}$ and $A_{m}^{\prime}$ the quotient rings $\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}$ and $\mathfrak{o}_{F^{\prime}}/\mathfrak{p}_{F^{\prime}}^{m}$ respectively. Recall that the isomorphism $\Lambda:A_{m}\xrightarrow{\sim}A_{m}^{\prime}$ takes the class of $\varpi$ to that of $\varpi^{\prime}$.
4.2.1. Unramified case
Suppose $E/F$ is unramified. In this case, the field $E$ is of the form $F[T]/(f)$, where $f$ is a monic irreducible polynomial with coefficients in $\mathfrak{o}_{F}$ such that its ${\rm mod}-\mathfrak{p}_{F}$ reduction, say $\overline{f}$, is irreducible. Moreover, the quotient ring $\mathfrak{o}_{E}/\mathfrak{p}_{E}^{m}$ is isomorphic to $A_{m}[T]/(\overline{f})$. Let $\overline{f}^{\prime}$ be the image of $\overline{f}$ under the ring isomorphism $\Lambda$. Then we have
$\mathfrak{o}_{E}/\mathfrak{p}_{E}^{m}\xrightarrow{\sim}A_{m}^{\prime}[T]/(\overline{f}^{\prime}).$
Let $f^{\prime}$ be an irreducible polynomial in $\mathfrak{o}_{F^{\prime}}[T]$ with its ${\rm mod}-\mathfrak{p}_{F}^{m}$ reduction $\overline{f}^{\prime}$. Let $E^{\prime}$ be the field extension $F^{\prime}[T]/(f^{\prime})$. Then $E^{\prime}$ is unramified over $F^{\prime}$ and there is a ring isomorphism
$\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{m}\simeq A_{m}^{\prime}[T]/(\overline{f}^{\prime}).$
Thus we have the ring isomorphism $\mathfrak{o}_{E}/\mathfrak{p}_{E}^{m}\xrightarrow{\sim}\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{m}$, which implies that the fields $E$ and $E^{\prime}$ are $m$-close.
4.2.2. Totally ramified case
Suppose $E$ is totally ramified over $F$. Here the field $E$ is of the form $F[T]/(T^{l}-\varpi)$ and its ring of integers $\mathfrak{o}_{E}$ is given by $\mathfrak{o}_{F}[T]/(T^{l}-\varpi)$. Recall that the class of $\varpi$ corresponds to the class of $\varpi^{\prime}$ via the isomorphism $\Lambda$. Now consider the field extension
$E^{\prime}:=F[T]/(T^{l}-\varpi^{\prime})$
Then $E^{\prime}$ is totally ramified over $F^{\prime}$ with the ring of integers $\mathfrak{o}_{E^{\prime}}=\mathfrak{o}_{F^{\prime}}[T]/(T^{l}-\varpi^{\prime})$. Let $\overline{\varpi}$ be the image of $\varpi$ under ${\rm mod}-\mathfrak{p}_{F}^{m}$. Then the natural surjection
$\mathfrak{o}_{F}[T]/(T^{l}-\varpi)\longrightarrow A_{m}[T]/(T^{l}-\overline{\varpi})$
factorzies through the ideal $\mathfrak{p}_{F}^{m}[T]/(T^{l}-\varpi)$, and it gives the ring isomorphism
$\mathfrak{o}_{E}/\mathfrak{p}_{E}^{lm}\xrightarrow{\sim}A_{m}[T]/(T^{l}-\overline{\varpi}).$
Similarly, there is a ring isomorphism
$\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{lm}\simeq A_{m}^{\prime}[T]/(T^{l}-\overline{\varpi^{\prime}}),$
where $\overline{\varpi^{\prime}}$ is the image of $\varpi^{\prime}$ under the ${\rm mod}-\mathfrak{p}_{F^{\prime}}^{m}$ reduction map. On the other hand, the map $\Lambda$ induces the ring isomorphism
$A_{m}[T]/(T^{l}-\overline{\varpi})\xrightarrow{\sim}A_{m}^{\prime}[T]/(T^{l}-\overline{\varpi^{\prime}}).$
This shows that the ring $\mathfrak{o}_{E}/\mathfrak{p}_{E}^{lm}$ is isomorphic to $\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{lm}$, and hence the fields $E$ and $E^{\prime}$ are $lm$-close. ∎
4.2.3.
For each $e\in\{1,l\}$, we fix a ring isomorphism $\Pi:\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}\rightarrow\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{em}$ as above, induced by $\Lambda$, so that the diagram
${\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Lambda}$${\mathfrak{o}_{F^{\prime}}/\mathfrak{p}_{F^{\prime}}^{m}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$${\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{em}}$
is commutative, where the vertical arrows are the natural inclusion maps. We choose uniformizer $\pi$ (resp. $\pi^{\prime}$) of the field $E$ (resp. $E^{\prime}$) such that the class of $\pi$ is mapped to the class of $\pi^{\prime}$ under the isomorphism $\Pi$. For any $\overline{a}\in\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}$ (or, $\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}$), we write $\overline{a}^{\prime}$ for the image $\Lambda(\overline{a})$ (or, $\Pi(\overline{a})$), whenever no confusion arises.
4.3. Compatibility with Galois action
Here, we continue to work with the set up of the preceeding subsection. We fix a generator $\sigma$ of the group ${\rm Gal}(E/F)$. Then $\sigma$ can be regarded as an automorphism of the group $G_{E}$, via its natural action on $G_{E}$. Consider the compact open subgroup
$K_{E}={\rm Ker}\big{(}\textbf{G}(\mathfrak{o}_{E})\rightarrow\textbf{G}(\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em})\big{)}.$
We observe that $K_{E}$ is $\sigma$-plain, and there is an induced action of $\langle\sigma\rangle$ on the Hecke algebra $\mathcal{H}(G_{E},K_{E})$, given by
$(\sigma.f)(x)=f(\sigma^{-1}x),$
for all $x\in G_{E}$ and $f\in\mathcal{H}(G_{E},K_{E})$. Similarly, there is an action of $\langle\sigma^{\prime}\rangle$ on the Hecke algebra $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})$. We choose a group isomorphism $\iota:{\rm Gal}(E/F)\rightarrow{\rm Gal}(E^{\prime}/F^{\prime})$, sending $\sigma$ to a generator, say $\sigma^{\prime}$, that makes commutative the diagram
${\mathfrak{o}_{E}/\mathfrak{p}^{em}_{E}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$$\scriptstyle{\sigma}$${\mathfrak{o}_{E}/\mathfrak{p}^{em}_{E}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$${\mathfrak{o}_{E^{\prime}}/\mathfrak{p}^{em}_{E^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\sigma^{\prime}}$${\mathfrak{o}_{E^{\prime}}/\mathfrak{p}^{em}_{E^{\prime}}}$
where the horizontal maps are the ring isomorphisms induced by $\sigma:E\rightarrow E$ and $\sigma^{\prime}:E^{\prime}\rightarrow E^{\prime}$. To be more precise, for any $a\in\mathfrak{o}_{E}$, we have the relation
$\overline{\sigma(a)}^{\prime}=\sigma(\overline{a})^{\prime}=\sigma^{\prime}(\overline{a}^{\prime})=\overline{\sigma^{\prime}(a^{\prime})},$ (4.3)
where $a^{\prime}\in\mathfrak{o}_{E^{\prime}}$ is the preimage of $\overline{a}^{\prime}$ under ${\rm mod}-\mathfrak{p}_{E^{\prime}}^{em}$. We end this section with the following proposition, which shows the compatibility of Kazhdan isomorphism with these Galois actions.
Proposition 4.3.
For all $f\in\mathcal{H}(G_{E},K_{E})$, we have
${\rm Kaz}_{em}^{E}(\sigma.f)=\sigma^{\prime}.{\rm Kaz}_{em}^{E}(f).$
Proof.
In view of the Cartan decomposition (4.2), it is enough to show that
${\rm Kaz}_{em}^{E}(\sigma.t_{a_{i}\pi_{\mu}a_{j}^{-1}})=\sigma^{\prime}.{\rm Kaz}_{em}^{E}(t_{a_{i}\pi_{\mu}a_{j}^{-1}}),$
for all $\mu\in X_{*}(\textbf{T})^{-}$ and $(a_{i},a_{j})\in T_{\mu}(E)$. Note that
$\sigma.t_{a_{i}\pi_{\mu}a_{j}^{-1}}=t_{\sigma(a_{i})\sigma(\pi_{\mu})\sigma(a_{j}^{-1})}.$
By Cartan decomposition (4.2), the double coset $K_{E}\sigma(a_{i})\sigma(\pi_{\mu})\sigma(a_{j}^{-1})K_{E}$ is of the form $K_{E}c_{i}\pi_{\mu}c_{j}^{-1}K_{E}$, where $(c_{i},c_{j})\in T_{\mu}(E)$. Then
${\rm Kaz}_{em}^{E}(\sigma.t_{a_{i}\pi_{\mu}a_{j}^{-1}})={\rm Kaz}_{em}^{E}\big{(}t_{\sigma(a_{i})\sigma(\pi_{\mu})\sigma(a_{j})^{-1}}\big{)}={\rm Kaz}_{em}^{E}\big{(}t_{c_{i}\pi_{\mu}c_{j}^{-1}}\big{)}=t_{c_{i}^{\prime}\pi_{\mu}^{\prime}c_{j}^{\prime-1}}.$ (4.4)
Consider the Galois conjugate $\sigma(\pi)=\pi t=\pi\sigma(s)$, where $t,s\in\mathfrak{o}_{E}^{\times}$. Under the relation $\sigma(\mu(x))=\mu(\sigma(x))$, for $x\in E^{\times}$, we observe that
$K_{E}c_{i}\pi_{\mu}c_{j}^{-1}K_{E}$ $=K_{E}\sigma(a_{i})\sigma(\pi_{\mu})\sigma(a_{j}^{-1})K_{E}$
$=K_{E}\sigma(a_{i})\pi_{\mu}\sigma(s_{\mu}a_{j}^{-1})K_{E}$
$=K_{E}\sigma(a_{i})\pi_{\mu}\sigma(d_{j}^{-1})K_{E},$
where $s_{\mu}=\mu(s)$, and $d_{j}=a_{j}s_{\mu}^{-1}\in\textbf{G}(\mathfrak{o}_{E})$. Then $\big{(}\bar{c_{i}}^{-1}\overline{\sigma(a_{i})},\overline{\sigma(d_{j}^{-1})}\bar{c_{j}}\big{)}\in\Gamma_{\mu}(E)$, and using the bijection $\Gamma_{\mu}(E)\rightarrow\Gamma_{\mu}(E^{\prime})$, we get
$\big{(}\bar{c_{i}}^{\prime-1}\overline{\sigma(a_{i})}^{\prime},\overline{\sigma(d_{j}^{-1})}^{\prime}\bar{c_{j}}^{\prime}\big{)}\in\Gamma_{\mu}(E^{\prime}).$
Recall that the image of $d_{j}$ under the reduction map $\textbf{G}(\mathfrak{o}_{E})\rightarrow\textbf{G}(\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em})$ is denoted by $\overline{d_{j}}$, and $\overline{d_{j}}^{\prime}$ is the image of $\overline{d_{j}}$ under the isomorphism $\textbf{G}(\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em})\xrightarrow{\sim}\textbf{G}(\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{em})$. Let $d_{j}^{\prime}\in\textbf{G}(\mathfrak{o}_{E^{\prime}})$ be a preimage of $\overline{d_{j}}^{\prime}$ under ${\rm mod}-\mathfrak{p}_{E^{\prime}}^{em}$. Then the commutative diagram
${\mathbb{G}_{m}(\mathfrak{o}_{E}/\mathfrak{p}^{em}_{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$$\scriptstyle{\mu}$${\textbf{T}(\mathfrak{o}_{E}/\mathfrak{p}^{em}_{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$${\mathbb{G}_{m}(\mathfrak{o}_{E^{\prime}}/\mathfrak{p}^{em}_{E^{\prime}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu}$${\textbf{T}(\mathfrak{o}_{E^{\prime}}/\mathfrak{p}^{em}_{E^{\prime}})}$
gives the relation
$\overline{d_{j}^{\prime}}=\overline{a_{j}^{\prime}}\overline{s_{\mu}}^{\prime-1}=\overline{a_{j}^{\prime}}(\overline{s^{\prime}_{\mu}})^{-1},$
where $s^{\prime}\in\mathfrak{o}_{E^{\prime}}$ is a preimage of $\overline{s}^{\prime}$ under ${\rm mod}-\mathfrak{p}_{E^{\prime}}^{em}$. On the other hand, the relation $\sigma(\pi)=\pi\sigma(s)$ together with (4.3) implies that
$\overline{\sigma^{\prime}(\pi^{\prime})}=\sigma^{\prime}(\overline{\pi}^{\prime})=\overline{\pi^{\prime}}\overline{\sigma^{\prime}(s^{\prime})}.$
For any $\mu\in X_{*}(\textbf{T})^{-}$, the above equality gives
$\overline{(\sigma^{\prime}(\pi^{\prime}))_{\mu}}=\overline{(\pi^{\prime}\sigma^{\prime}(s^{\prime}))_{\mu}}.$
This implies that $(\sigma^{\prime}(\pi^{\prime}))_{\mu}^{-1}(\pi^{\prime}\sigma^{\prime}(s^{\prime}))_{\mu}\in K_{E^{\prime}}$. Now, using normality of $K_{E^{\prime}}$ in $\textbf{G}(\mathfrak{o}_{E^{\prime}})$ and the relation $\sigma(\mu(x))=\mu(\sigma(x))$, for $x\in E^{\times}$, we get
$K_{E^{\prime}}c_{i}^{\prime}\pi_{\mu}^{\prime}c_{j}^{\prime-1}K_{E^{\prime}}$ $=K_{E^{\prime}}\sigma^{\prime}(a_{i}^{\prime})\pi_{\mu}^{\prime}\sigma^{\prime}(d_{j}^{\prime-1})K_{E^{\prime}}$
$=K_{E^{\prime}}\sigma^{\prime}(a_{i}^{\prime})\sigma^{\prime}(\pi_{\mu}^{\prime})\sigma^{\prime}(a_{j}^{\prime-1})K_{E^{\prime}}.$
This shows that
$t_{c_{i}^{\prime}\pi_{\mu}^{\prime}c_{j}^{\prime-1}}=t_{\sigma^{\prime}(a_{i}^{\prime})\sigma^{\prime}(\pi_{\mu}^{\prime})\sigma^{\prime}(a_{j}^{\prime-1})}=\sigma^{\prime}.t_{a_{i}^{\prime}\pi_{\mu}^{\prime}a_{j}^{\prime-1}}.$
Finally, it follows from (4.4) and the above relation that
${\rm Kaz}_{em}^{E}(\sigma.t_{a_{i}\pi_{\mu}a_{j}^{-1}})=\sigma^{\prime}.{\rm Kaz}_{em}^{E}(t_{a_{i}\pi_{\mu}a_{j}^{-1}}).$
5. Compatibility of Brauer homomorphism and Kazhdan isomorphism
In this section, we prove Theorem 1.1. Let us first recall the notations and terminologies from the prceeding sections.
5.1.
Fix a positive integer $m$. Let $F$ and $F^{\prime}$ be the non-Archimedean local fields such that $F$ is $m$-close to $F^{\prime}$. Let $E$ be a finite Galois extension of $F$ of prime degree $l$ with $l\neq p$, where $p$ is the residue characteristic of both $F$ and $F^{\prime}$. According to Lemma 4.2, there exists a finite Galois extension $E^{\prime}$ of $F^{\prime}$ of degree $l$ and $E^{\prime}$ is $em$-close to $E$, where $e\in\{1,l\}$.
Let G be a split connected reductive group defined over $\mathbb{Z}$. It follows from [KP23, Proposition 12.9.4] that the fixed point subgroup $K_{E}^{\sigma}$ (resp. $K_{E^{\prime}}^{\sigma^{\prime}}$) is the $m$-th congruence subgroup of $\textbf{G}(\mathfrak{o}_{F})$ (resp. $\textbf{G}(\mathfrak{o}_{F^{\prime}})$) i.e, $K_{E}^{\sigma}=K_{F}$ and $K_{E^{\prime}}^{\sigma^{\prime}}=K_{F^{\prime}}$, where
$K_{F}={\rm Ker}\big{(}\textbf{G}(\mathfrak{o}_{F})\rightarrow\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})\big{)}$
and
$K_{F^{\prime}}={\rm Ker}\big{(}\textbf{G}(\mathfrak{o}_{F^{\prime}})\rightarrow\textbf{G}(\mathfrak{o}_{F^{\prime}}/\mathfrak{p}_{F^{\prime}}^{m})\big{)}.$
Proposition 4.3 then gives an isomorphism between $\mathcal{H}(G_{E},K_{E})^{\sigma}$ and $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}$, induced by ${\rm Kaz}_{em}^{E}$. On the other hand, we have the Brauer homomorphism $\overline{\rm Br}$ (resp. $\overline{\rm Br}^{\prime}$) from the space $\mathcal{H}(G_{E},K_{E})^{\sigma}$ (resp. $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}$) to the Hecke algebra $\mathcal{H}(G_{F},K_{F})$ (resp. $\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$). With these set up, we aim to prove
Theorem 5.1.
Let $F,F^{\prime},E$ and $E^{\prime}$ be as above. Then for any split connected reductive group G defined over $\mathbb{Z}$, we have
${\rm Kaz}_{m}^{F}\circ\overline{\rm Br}=\overline{\rm Br}^{\prime}\circ{\rm Kaz}_{em}^{E}.$
Proof.
The proof is divided into two parts. In the first part, we consider the extension $E/F$ to be unramified. The second part deals with totally ramified case. To start with, let us recall that an arbitrary $h\in\mathcal{H}(G_{E},K_{E})^{\sigma}$ can be written as
$h=\sum_{\mu\in X_{*}(\textbf{T})^{-}}\sum_{(a_{i},a_{j})\in T_{\mu}(E)}\alpha_{\mu}(a_{i},a_{j})t_{a_{i}\pi_{\mu}a_{j}^{-1}},$
where each $\alpha_{\mu}(a_{i},a_{j})\in k$. Then
$({\rm Kaz}_{m}^{F}\circ{\rm\overline{Br}})(h)=\sum_{(a_{i},a_{j})\in T_{\mu}(E)}\alpha_{\mu}(a_{i},a_{j})({\rm Kaz}_{m}^{F}\circ{\rm\overline{Br}})(t_{a_{i}\pi_{\mu}a_{j}^{-1}})$ (5.1)
and
$({\rm\overline{Br}}^{\prime}\circ{\rm Kaz}_{em}^{E})(h)=\sum_{\mu\in X_{*}(\textbf{T})^{-}}\sum_{(a_{i},a_{j})\in T_{\mu}(E^{\prime})}\alpha_{\mu}(a_{i},a_{j}){\rm\overline{Br}}^{\prime}(t_{a_{i}^{\prime}\pi_{\mu}^{\prime}a_{j}^{\prime-1}}).$ (5.2)
For each $\mu\in X_{*}(\textbf{T})^{-}$ and $(a_{i},a_{j})\in T_{\mu}(E)$, we have ${\rm\overline{Br}}(t_{a_{i}\pi_{\mu}a_{j}^{-1}})=1_{G_{F}\cap K_{E}a_{i}\pi_{\mu}a_{j}^{-1}K_{E}}$. Then Cartan decomposition (4.2) gives
$K_{E}a_{i}\pi_{\mu}a_{j}^{-1}K_{E}\cap G_{F}=\coprod_{\lambda\in X_{*}(\textbf{T})^{-}}\coprod_{(b_{i},b_{j})\in T_{\lambda}(F)}K_{E}a_{i}\pi_{\mu}a_{j}^{-1}K_{E}\cap K_{F}b_{i}\varpi_{\lambda}b_{j}^{-1}K_{F}.$ (5.3)
Similarly,
$K_{E^{\prime}}a_{i}^{\prime}\pi_{\mu}^{\prime}a_{j}^{\prime-1}K_{E^{\prime}}\cap G_{F^{\prime}}=\coprod_{\nu\in X_{*}(\textbf{T})^{-}}\coprod_{(d_{i}^{\prime},d_{j}^{\prime})\in T_{\nu}(F^{\prime})}K_{E^{\prime}}a_{i}^{\prime}\pi^{\prime}_{\mu}a_{j}^{\prime-1}K_{E^{\prime}}\cap K_{F^{\prime}}d_{i}^{\prime}\varpi^{\prime}_{\nu}d_{j}^{\prime-1}K_{F^{\prime}}.$ (5.4)
5.1.1. Unramified case
Suppose $E$ is unramified over $F$. Then the field $E^{\prime}$ is also unramified over $F^{\prime}$ and it is $m$-close to $E$ (Lemma 4.2). In this case, $\pi=\varpi$ and $\pi^{\prime}=\varpi^{\prime}$. For any $\lambda\in X_{*}(\textbf{T})^{-}$ and $(b_{i},b_{j})\in T_{\lambda}(F)$, it follows from the decompositions (5.3) and (5.4) that
$b_{i}\varpi_{\lambda}b_{j}^{-1}\in K_{E}a_{i}\varpi_{\mu}a_{j}^{-1}K_{E}$
if and only if
$\varpi_{\lambda}\in K_{E}b_{i}^{-1}a_{i}\varpi_{\mu}a_{j}^{-1}b_{j}K_{E},$
and this will happen only if $\lambda=\mu$. Then
$\varpi_{\mu}\in K_{E}b_{i}^{-1}a_{i}\varpi_{\mu}a_{j}^{-1}b_{j}K_{E}$
if and only if
$(b_{i}^{-1}a_{i},b_{j}^{-1}a_{j})\in\Gamma_{\mu}(E).$
Using the bijection $\Gamma_{\mu}(E)\rightarrow\Gamma_{\mu}(E^{\prime})$ $(\overline{a}\mapsto\overline{a}^{\prime})$, we have
$(\overline{b_{i}}^{-1}\overline{a_{i}},\overline{b_{j}}^{-1}\overline{a_{j}})\in\Gamma_{\mu}(E)$
if and only if
$(\overline{b_{i}}^{\prime-1}\overline{a_{i}}^{\prime},\overline{b_{j}}^{\prime-1}\overline{a_{j}}^{\prime})=(\overline{b_{i}^{\prime}}^{-1}\overline{a_{i}^{\prime}},\overline{b_{j}^{\prime}}^{-1}\overline{a_{j}^{\prime}})\in\Gamma_{\mu}(E^{\prime}),$
where $(b_{i}^{\prime},b_{j}^{\prime})\in T_{\mu}(F^{\prime})$. Therefore, we have
$b_{i}\varpi_{\mu}b_{j}^{-1}\in K_{E}a_{i}\varpi_{\mu}a_{j}^{-1}K_{E}\iff b_{i}^{\prime}\varpi_{\mu}^{\prime}b_{j}^{\prime-1}\in K_{E^{\prime}}a_{i}^{\prime}\varpi_{\mu}^{\prime}a_{j}^{\prime-1}K_{E^{\prime}}$ (5.5)
In view of (5.5), it follows from (5.1) and (5.2) that
$({\rm Kaz}_{m}^{F}\circ{\rm\overline{Br}})(h)=({\rm\overline{Br}}^{\prime}\circ{\rm Kaz}_{m}^{E})(h).$ (5.6)
5.1.2. Totally Ramified case
Here, we assume that the extension $E/F$ is totally ramified. Then Lemma 4.2 ensures that the extension $E^{\prime}/F^{\prime}$ is also totally ramified and $E^{\prime}$ is $lm$-close to $E$, where $l=[E:F]=[E^{\prime}:F^{\prime}]$. Note that
$K_{E}={\rm Ker}\big{(}\textbf{G}(\mathfrak{o}_{E})\xrightarrow{{\rm mod}-\mathfrak{p}_{E}^{lm}}\textbf{G}(\mathfrak{o}_{E}/\mathfrak{p}_{E}^{lm})\big{)}$
and
$K_{F}={\rm Ker}\big{(}\textbf{G}(\mathfrak{o}_{F})\xrightarrow{\text{mod}-\mathfrak{p}_{F}^{m}}\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})\big{)}.$
Let $\xi\in X_{*}(\textbf{T})^{-}$ and $(c_{i},c_{j})\in T_{\xi}(F)$. In view of the decompositions (5.3) and (5.4), we have
$c_{i}\varpi_{\xi}c_{j}^{-1}\in K_{E}a_{i}\pi_{\mu}a_{j}^{-1}K_{E}$
if and only if
$\varpi_{\xi}\in K_{E}c_{i}^{-1}a_{i}\pi_{\mu}a_{j}^{-1}c_{j}K_{E},$
and it is possible only if $\xi^{l}=\mu$. Therefore, it follows that
$\pi_{\xi^{l}}\in K_{E}c_{i}^{-1}a_{i}\pi_{\xi^{l}}a_{j}^{-1}c_{j}K_{E}$
if and only if
$(\overline{c_{i}}^{-1}\overline{a_{i}},\overline{a_{j}}^{-1}\overline{c_{j}})\in\Gamma_{\xi^{l}}(E).$
Under the bijection $\Gamma_{\xi^{l}}(E)\rightarrow\Gamma_{\xi^{l}}(E^{\prime})$, we have
$(\overline{c_{i}}^{-1}\overline{a_{i}},\overline{a_{j}}^{-1}\overline{c_{j}})\in\Gamma_{\xi^{l}}(E^{\prime})$
if and only if
$(\overline{c_{i}}^{\prime-1}\overline{a_{i}}^{\prime},\overline{c_{j}}^{\prime-1}\overline{a_{j}}^{\prime})=(\overline{c_{i}^{\prime}}^{-1}\overline{a_{i}^{\prime}},\overline{c_{j}^{\prime}}^{-1}\overline{a_{j}^{\prime}})\in\Gamma_{\xi^{l}}(E^{\prime}),$
where $(c_{i}^{\prime},c_{j}^{\prime})\in T_{\xi^{l}}(F^{\prime})$. Thus we have
$c_{i}\varpi_{\xi}c_{j}^{-1}\in K_{E}a_{i}\pi_{\mu}a_{j}^{-1}K_{E}\iff c_{i}^{\prime}\varpi^{\prime}_{\xi}c_{j}^{\prime-1}\in K_{E^{\prime}}a_{i}^{\prime}\pi^{\prime}_{\mu}a_{j}^{\prime-1}K_{E^{\prime}}.$ (5.7)
From the relations (5.1), (5.2) and (5.7), we get
$({\rm Kaz}_{m}^{F}\circ{\rm\overline{Br}})(h)=({\rm\overline{Br}}^{\prime}\circ{\rm Kaz}_{lm}^{E})(h).$
This completes the proof. ∎
6. Linkage over close local fields
In this section, an application of Theorem 5.1 is given in the context of linkage. To be more precise, we show that the linkage is compatible with Kazhdan isomorphism. To prove this, we first give a formulation of the compatibility of linkage with Brauer homomorphism, as described in [TV16, Section 6.2], and then we use Theorem 5.1 to get the required compatibility. To start with, we first recall the Tate cohomology and the concept of linkage. For a reference, see [TV16, Section 3 and 6].
6.1. Tate cohomology
Let $\Gamma$ be a cyclic group of order $n$ with generator $\gamma$. Let $M$ be an abelian group with an action of $\Gamma$. Let $T_{\gamma}$ be the automorphism of $M$ defined by
$T_{\gamma}(m)=\gamma.m$, for $\gamma\in\Gamma,m\in M$.
Let $N={\rm id}+T_{\gamma}+T_{\gamma^{2}}+\dots+T_{\gamma^{n-1}}$ be the norm operator. The Tate cohomology groups $\widehat{H}^{0}(M)$ and $\widehat{H}^{1}(M)$ are defined as
$\widehat{H}^{0}(M):={\rm ker}({\rm id}-T_{\gamma})/{\rm Im}(N)$,
$\widehat{H}^{1}(M):={\rm ker}(N)/{\rm Im}({\rm id}-T_{\gamma})$.
6.1.1. Tate Cohomology of sheaves on $\ell$-spaces
Let $X$ be an $\ell$-space (i.e., Hausdorff, locally compact and totally disconnected) with an action of a finite group $\langle\gamma\rangle$ of prime order $l$. Let $\mathcal{F}$ be a sheaf of $k$-modules on $X$. If $\mathcal{F}$ is $\gamma$-equivariant, then $\gamma$ can be regarded as a map of restricted sheaves $\mathcal{F}|_{X^{\gamma}}\rightarrow\mathcal{F}|_{X^{\gamma}}$ and the Tate cohomology is defined as
$\widehat{H}^{0}(\mathcal{F}|_{X^{\gamma}}):={\rm ker}(1-\gamma)/{\rm Im}(N)$,
$\widehat{H}^{1}(\mathcal{F}|_{X^{\gamma}}):={\rm ker}(N)/{\rm Im}(1-\gamma)$.
For each $i\in\{0,1\}$, the object $\widehat{H}^{i}(\mathcal{F}|_{X^{\gamma}})$ is a sheaf on $X^{\gamma}$.
6.2. Linkage
Let $F$ be a non-archimedean local field and let $G$ be the $F$-points of a connected reductive group defined over $F$. A representation $(\rho,V)$ of $G$ is called smooth if for every vector $v\in V$, the $G$-stabilizer of $v$ is an open subgroup of $G$. Moreover, such a representation $\rho$ is called admissible if for each compact open subgroup $K$ of $G$, the space of $K$-fixed vectors $V^{K}$ (sometimes denoted by $\rho^{K}$) is finite-dimensional over $k$. In this subsection, all the representations are assumed to be admissible and the representation spaces are defined over $k$. We denote by $\mathcal{R}(G)$ the category of smooth admissible $k$-representations of $G$.
6.2.1. Frobenius Twist
Let $(\rho,V)$ be a representation of $G$. Consider the vector space $V^{(l)}$, where the underlying additive group structure of $V^{(l)}$ is same as that of $V$ but the scalar action $\star$ on $V^{(l)}$ is given by
$c\star v=c^{\frac{1}{l}}.v$ , for all $c\in k,v\in V$.
Then the action of $G$ on $V$ induces a representation $\rho^{(l)}$ of $G$ on $V^{(l)}$. The representation $(\rho^{(l)},V^{(l)})$ is called the Frobenius twist of the representation $(\rho,V)$.
6.2.2.
Let $\sigma$ be an automorphism of $G$ of prime order $l$. An irreducible representation $(\Xi,W)$ of $G$ is called $\sigma$-fixed if $\Xi$ is isomorphic to $\Xi\circ\sigma$. For such representation $\Xi$, there is an unique action of $\sigma$ on the space $W$, which is compatible with the action of $\sigma$ on $G$ ([TV16, Proposition 6.1]). Then we have the Tate cohomology space $\widehat{H}^{i}(W)$, defined as $k$-representation of $G^{\sigma}$. We denote this representation by $\widehat{H}^{i}(\Xi)$.
Definition 6.1.
An irreducible representation $\rho$ of $G^{\sigma}$ is linked with the representation $\Xi$ if the Frobenius-twist $\rho^{(l)}$ occurs as a Jordan-Holder constituent of $\widehat{H}^{0}(\Xi)$ or $\widehat{H}^{1}(\Xi)$.
6.2.3.
It was conjectured by Venkatesh-Treumann ([TV16, Section 6.3]) that $\widehat{H}^{i}(\Xi)$ is of finite length as a representation of $G^{\sigma}$. We give a proof of this conjecture for the group ${\rm GL}_{n}$, under the following set up.
[…]
We start with a non-archimedean local field $L$. Let $P_{n}(L)$ and $U_{n}(L)$ be the subgroups of ${\rm GL}_{n}(L)$, given by
$P_{n}(L)=\Big{\{}\begin{pmatrix}A&B\\ 0&1\end{pmatrix}:A\in{\rm GL}_{n-1}(L),B\in L^{n-1}\Big{\}}$
and
$U_{n}(L)=\Big{\{}\begin{pmatrix}1&C\\ 0&1\end{pmatrix}:C\in L^{n-1}\Big{\}}$
respectively. We denote by $X_{L}$ the coset space $P_{n}(L)/U_{n}(L)$. Let $\theta_{n}$ be the character of $U_{n}(L)$, defined as
$\theta_{n}((u)_{i,j}):=\psi_{L}(u_{n-1,n})$, for $(u)_{i,j}\in U_{n}(L)$.
We have four fundamental functors:
$\Psi^{-}:\mathcal{R}(P_{n}(L))\rightarrow\mathcal{R}({\rm GL}_{n-1}(L)),\Psi^{+}:\mathcal{R}({\rm GL}_{n-1}(L))\rightarrow\mathcal{R}(P_{n}(L))$
$\Phi^{-}:\mathcal{R}(P_{n}(L))\rightarrow\mathcal{R}(P_{n-1}(L)),\Phi^{+}:\mathcal{R}(P_{n-1}(L))\rightarrow\mathcal{R}(P_{n}(L)),$
given by $\Psi^{-}=r_{U_{n}(L),{\rm id}}$, $\Phi^{-}=r_{U_{n}(L),\theta_{n}}$, $\Psi^{+}=i_{U_{n}(L),{\rm id}}$, and $\Phi^{+}=i_{U_{n}(L),\theta_{n}}$.
Let $\tau$ be a smooth representation of $P_{n}(L)$. The $m$-th derivative of $\tau$, denoted by $\tau^{(m)}$, is constructed as the representation $\Psi^{-}(\Phi^{-})^{m-1}(\tau)$ of ${\rm GL}_{n-m}(L)$. Moreover, we get a functorial filtration on $\tau$:
$0\subseteq\tau_{n}\subseteq\tau_{n-1}\subseteq\dots\subseteq\tau_{2}\subseteq\tau_{1}=\tau,$
where $\tau_{m}=(\Phi^{+})^{m-1}(\Phi^{-})^{m-1}(\tau)$ and $\tau_{m}/\tau_{m+1}=(\Phi^{+})^{m-1}(\Psi^{+})(\tau^{(m)})$. Using these, we now deduce:
Proposition 6.2.
Let $\Xi$ be an irreducible $\sigma$-fixed representation of ${\rm GL}_{n}(E)$. Let $T_{\sigma}:\Xi\rightarrow\Xi\circ\sigma$ be an isomorphism with $T_{\sigma}^{l}={\rm id}$. Then the Tate cohomology $\widehat{H}^{i}(\Xi)$ with respect to the operator $T_{\sigma}$ is a finite length representation of ${\rm GL}_{n}(F)$.
Proof.
We use induction on the integer $n$. Assume that the proposition is true for all positive integers less than $n$. Since $\psi_{E}(\sigma(x))=\psi_{E}(x)$, for $x\in E$, we get the isomorphism
$\Phi^{-}(\Xi\circ\sigma)\simeq\Phi^{-}(\Xi)\circ\sigma,$ (6.1)
as representation of $P_{n-1}(E)$. Similarly, for any representation $\tau\in\mathcal{R}(P_{n-1}(E))$, we have the $P_{n}(E)$-equivariant isomorphism
$\Phi^{+}(\tau\circ\sigma)\simeq\Phi^{+}(\tau)\circ\sigma.$ (6.2)
The map $T_{\sigma}$ together with (6.1) and (6.2) induces an isomorphism between $\Xi^{(m)}$ and $(\Xi^{(m)})\circ\sigma$, and also between the representations $\Xi_{m}$ and $\Xi_{m}\circ\sigma$ for $m\geq 1$. Using induction hypothesis, we get that the representation $\widehat{H}^{i}(\Xi^{(m)})$ is a finite length representation of ${\rm GL}_{n-m}(F)$. Since $P_{n}(E)$ is stable under the action of $\sigma$, for any $n\geq 1$, we get that there is a bijection between the sets $X_{E}^{\sigma}$ and $X_{F}$. Then according to [TV16, Proposition 3.3.1], we have
$\widehat{H}^{i}(\Xi_{m}/\Xi_{m+1})=(\Phi^{+})^{m-1}(\Psi^{+})\big{(}\widehat{H}^{i}(\Xi^{(m)})\big{)}.$
Hence, for all $m\geq 1$, the representation $\widehat{H}^{i}(\Xi_{m}/\Xi_{m+1})$ is finite length representation of $P_{n}(F)$. Now, using spectral sequence induced by canonical filtration of ${\rm Res}_{P_{n}(E)}(\Xi)$, we get that $\widehat{H}^{i}(\Xi)$ is finite length representation of $P_{n}(F)$, and this implies that $\widehat{H}^{i}(\Xi)$ is finite length representation of ${\rm GL}_{n}(F)$. ∎
It is reasonable to expect that Proposition 6.2 holds true for any general connected reductive algebraic group G over $F$. In the next section, we give a formulation of the compatibility of Brauer homomorphism with linkage, under the finite length assumption of the Tate cohomology $\widehat{H}^{i}(\Xi)$ for any irreducible $\sigma$-fixed representation $\Xi$ of $G=\textbf{G}(F)$. Then we show that this formulaton is compatible with Kazhdan isomorphism.
6.3. Compatibility of Brauer homomorphism with linkage
Let $\Xi$ be an irreducible $\sigma$-fixed representation of $G$, and let $T:\Xi\rightarrow\Xi\circ\sigma$ be an isomorphism such that $T^{l}={\rm id}$. For any $\sigma$-plain compact open subgroup $K$ of $G$, the inclusion map $\Xi^{K}\hookrightarrow\Xi$ is $\sigma$-invariant, and it induces a map $\widehat{H}^{i}(\Xi^{K})\rightarrow\widehat{H}^{i}(\Xi)$, taking values in the $\mathcal{H}(G^{\sigma},K^{\sigma})$-module $\widehat{H}^{i}(\Xi)^{K^{\sigma}}$. It is shown in [TV16, Section 6.2] that for all $h\in\mathcal{H}(G,K)^{\sigma}$, the Brauer homomorphism $\overline{\rm Br}$ induces the commutative diagram
${\widehat{H}^{i}(\Xi^{K})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widehat{H}^{i}(h)}$${\widehat{H}^{i}(\Xi)^{K^{\sigma}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\rm\overline{Br}}(h)}$${\widehat{H}^{i}(\Xi^{K})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{i}(\Xi)^{K^{\sigma}}}$
and this induces the compatibility of Brauer homomorphism with linkage in the sense that, for a representation $\rho$ of $G^{\sigma}$ to be linked with $\Xi$, it is equivalent to say that there exists a composition factor $N$ of the $\mathcal{H}(G,K)^{\sigma}$-module $\widehat{H}^{i}(\Xi^{K})$ such that
$\mathcal{H}(G^{\sigma},K^{\sigma})\otimes_{\mathcal{H}(G,K)^{\sigma}}N\simeq(\rho^{(l)})^{K^{\sigma}}.$ (6.3)
6.3.1. Compatibility of Kazhdan isomorphism with linkage
We first recall the notations and set up under which we study this compatibility. Let $F$ and $F^{\prime}$ be the non-Archimedean $m$-close local fields with same residue characteristic $p$. Let $E$ be the finite Galois extension of $F$ of prime degree $l$ with $l\neq p$. By Lemma 4.2, there exists a finite Galois extension $E^{\prime}/F^{\prime}$ with same degree of extension $l$. As before, we fix the generators $\sigma$ and $\sigma^{\prime}$ of the Galois groups ${\rm Gal}(E/F)$ and ${\rm Gal}(E^{\prime}/F^{\prime})$ respectively.
Let G be a split connected reductive group defined over $\mathbb{Z}$. Let $G_{E}$ (resp. $G_{F}$) be the $E$ (resp. $F$)-points of G. Let $K_{E}$ and $K_{F}$ be the $m$-th congruence subgroups of $G_{E}$ and $G_{F}$ respectively (see Section 4). Then we have
${\rm Kaz}_{m}^{F}\circ{\rm\overline{Br}}={\rm\overline{Br}^{\prime}}\circ{\rm Kaz}_{em}^{E},$ (6.4)
where $e\in\{1,l\}$. It is shown in [Gan15, Section 2.3] that there are bijections
$\big{\{}\rho_{E}\in{\rm Irr}(G_{E}):\rho_{E}^{K_{E}}\neq 0\big{\}}\longrightarrow\big{\{}\rho_{E^{\prime}}\in{\rm Irr}(G_{E^{\prime}}):\rho_{E^{\prime}}^{K_{E^{\prime}}}\neq 0\big{\}}$ (6.5)
and
$\big{\{}\rho_{F}\in{\rm Irr}(G_{F}):\rho_{F}^{K_{F}}\neq 0\big{\}}\longrightarrow\big{\{}\rho_{F^{\prime}}\in{\rm Irr}(G_{F^{\prime}}):\rho_{F^{\prime}}^{K_{F^{\prime}}}\neq 0\big{\}},$ (6.6)
induced by the Kazhdan isomorphisms ${\rm Kaz}_{em}^{E}$ and ${\rm Kaz}_{m}^{F}$ respectively. With these notations and terminologies, we now prove
Theorem 6.3.
Let $(\rho_{E},V_{E})$ be an irreducible $\sigma$-fixed representation of $G_{E}$ with $V_{E}^{K_{E}}\neq 0$, and let $(\rho_{F},V_{F})$ be an irreducible representation of $G_{F}$ with $V_{F}^{K_{F}}\neq 0$. Let $(\rho_{E^{\prime}},V_{E^{\prime}})$ and $(\rho_{F^{\prime}},V_{F^{\prime}})$ be the corresponding objects under the bijections (6.5) and (6.6). Then $\rho_{F}$ is linked with $\rho_{E}$ if and only if $\rho_{F^{\prime}}$ is linked with $\rho_{E^{\prime}}$.
Proof.
The proof is divided into two parts. In the first part, we show that the irreducible representation $\rho_{E^{\prime}}$, obtained from $\rho_{E}$ under the bijection (6.5), is $\sigma^{\prime}$-fixed. The second part shows that the linkage between $\rho_{F}$ and $\rho_{E}$ implies the linkage between $\rho_{F^{\prime}}$ and $\rho_{E^{\prime}}$, and vice versa.
6.3.2.
First recall that the space $V_{E}^{K_{E}}$ is an irreducible $\mathcal{H}(G_{E},K_{E})$-module, and it can be regarded as $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})$-module via Kazhdan isomorphism. Consider the $\mathcal{H}(G_{E^{\prime}})$ (For definition, see [Vig96, Chapter 1])-module
$N^{\prime}=\mathcal{H}(G_{E^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})}V_{E}^{K_{E}}.$
Note that the action of $\langle\sigma^{\prime}\rangle$ on $\mathcal{H}(G_{E^{\prime}})$ induces an action of $\langle\sigma^{\prime}\rangle$ on $N^{\prime}$, given by
$\sigma^{\prime}.(f^{\prime}\otimes v)=(\sigma^{\prime}.f^{\prime})\otimes v,$
for $f^{\prime}\in\mathcal{H}(G_{E^{\prime}})$ and $v\in V_{E}^{k_{E}}$. Following the construction of [Vig96, Section 4, Chapter 1], the $\mathcal{H}(G_{E^{\prime}})$-module $N^{\prime}$ gives the irreducible representation $\rho_{E^{\prime}}$, defined as
$\rho_{E^{\prime}}(g^{\prime})(x^{\prime})=1_{g^{\prime}K_{E^{\prime}}}*x^{\prime},$
for $g^{\prime}\in G_{E^{\prime}},x^{\prime}\in V_{E^{\prime}}$. Here, *** denotes the action of the Hecke algebra $\mathcal{H}(G_{E^{\prime}})$ on the space $V_{E^{\prime}}$ and $1_{g^{\prime}K_{E^{\prime}}}$ is the characteristic function of the left coset $g^{\prime}K_{E^{\prime}}$. Since $\rho_{E}$ is $\sigma$-fixed, there exists an isomorphism $T_{\sigma}:\rho_{E}\rightarrow\rho_{E}\circ\sigma$. Then the map $T_{\sigma^{\prime}}:N^{\prime}\rightarrow N^{\prime}$, defined by
$T_{\sigma^{\prime}}(f^{\prime}\otimes v):=(\sigma^{\prime}.f^{\prime})\otimes T_{\sigma}(v),$
induces an isomorphism $\rho_{E^{\prime}}\xrightarrow{\sim}\rho_{E^{\prime}}\circ\sigma^{\prime}$, and this shows that the representation $\rho_{E^{\prime}}$ is $\sigma^{\prime}$-fixed.
6.3.3.
Now, suppose that the representation $\rho_{F}$ is linked with $\rho_{E}$. Then the $\mathcal{H}(G_{F},K_{F})$-module $(V_{F}^{(l)})^{K_{F}}$ is a composition factor of $\widehat{H}^{i}(V_{E})^{K_{F}}$, for some $i\in\{0,1\}$. If we regard these as $\mathcal{H}(G_{E},K_{E})^{\sigma}$-modules via the Brauer homomorphism $\overline{\rm Br}$, then the formulation (6.3) gives a composition factor $M$ of $\widehat{H}^{i}(V_{E}^{K_{E}})$ such that
$\mathcal{H}(G_{F},K_{F})\otimes_{\mathcal{H}(G_{E},K_{E})^{\sigma}}M\simeq(V_{F}^{(l)})^{K_{F}}.$ (6.7)
Note that the isomorphism $V_{E}^{K_{E}}\rightarrow V_{E^{\prime}}^{K_{E^{\prime}}}$ of $\mathcal{H}(G_{E},K_{E})$-modules induces a $\mathcal{H}(G_{E},K_{E})^{\sigma}$-module isomorphism $\Phi:\widehat{H}^{i}(V_{E}^{K_{E}})\rightarrow\widehat{H}^{i}(V_{E^{\prime}}^{K_{E^{\prime}}})$, which is also a $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}$-module isomorphism via the Kazhdan map ${\rm Kaz}^{E}_{em}$.
Let $M^{\prime}$ be the image of $M$ under the map $\Phi$. Then $M^{\prime}$ is a composition factor of $\widehat{H}^{i}(V_{E^{\prime}}^{K_{E^{\prime}}})$, and the $\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$-module
$\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}}M^{\prime}$
is a composition factor of $\widehat{H}^{i}(V_{E^{\prime}})^{K_{F^{\prime}}}$. We also have an isomorphism $\varphi:V_{F}^{K_{F}}\rightarrow V_{F^{\prime}}^{K_{F^{\prime}}}$, which is equivariant under the action of $\mathcal{H}(G_{F},K_{F})$ and hence $\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$ via the Kazhdan isomorphism ${\rm Kaz}_{m}^{F}$. Using the isomorphisms $\Phi$, $\varphi$ and the relation (6.4), it follows from (6.7) that
$\mathcal{H}(G_{F,},K_{F^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}}M^{\prime}\simeq(V_{F^{\prime}}^{(l)})^{K_{F^{\prime}}}.$
This shows that the representation $\rho_{F^{\prime}}$ is linked with $\rho_{E^{\prime}}$. A similar argument also establishes the other implication, namely that the representation $\rho_{F}$ is linked with $\rho_{E}$ if $\rho_{F^{\prime}}$ is linked with $\rho_{E^{\prime}}$. This completes the proof. ∎
Acknowledgements. The author is grateful to Santosh Nadimpalli for communicating Proposition 6.2, and for his helpful comments during the work.